# Understanding Ranking LLMs: A Mechanistic Analysis for Information Retrieval

### v2

@@ 4.2. Comparing Ranking LLMs
We next explore whether different ranking LLMs encode similar or distinct Information Retrieval features. To investigate this, we extend our analysis beyond the RankLlama2-7b model and compare it to the two additional fine-tuned LLMs, RankLlama3-8b and RankPythia-6.9b, as well as the RankLlama2-13b model. We fine-tune these models under an identical passage reranking framework, with an identical dataset and loss function, to ensure a fair and consistent comparison.
@@
All four ranking LLMs demonstrate broadly comparable patterns for most features, particularly for the Min metrics (e.g., Min of TF, Min of TF*IDF), where they uniformly show top scores (all 1.0). They also appear closely aligned on coverage-based features—such as Covered Query Term Number and Ratio—with each scoring in the high 0.95–0.98 range. By contrast, the greatest differences emerge in TF-based statistics like Max of Term Frequency and Variance of TF/TF*IDF, where RankPythia tends to exhibit slightly higher peaks (e.g., a larger $R^{2}$ for Max of TF) than the RankLLaMA models. Additionally, all LLMs show near zero performance in encoding Sum of Stream Length Normalized TF, suggesting similarity in how these models handle term-frequency weighting. Overall, the four models encode IR features in a largely similar manner, indicating that our probing findings are generalizable and not LLM specific.
@@ 4.3. Out-of-distribution queries and documents
We compared the probing performance of in-distribution and out-of-distribution query-document pairs in RankLlama3-8b and RankPythia-6.9b, observing similar performance for both types of inputs, consistent with the results observed for RankLlama2-7b in Figure 5a. The graphs depicting these probing results are not included due to space constraints.
@@ 4.4. Feature Groups within RankLlama
In the previous subsection, we found that several MSLR features (covered QT Number, covered QT Ratio, mean(TF/L), min TF, min of TF*IDF, Min(TF/L) and variance of TF*IDF) appear to be modeled within LLM activations, especially within the later layers. These features also seem to be a fair choice to model relevance based on intuition. However, these features might not be present individually but in combination with one another or individually with different exponents. To test for this possibility, we probe for combinations of these features within different layers of the LLM and find various combinations of covered query term ratio, mean of stream length normalized term frequency and variance of tf*idf. We find strong indication for their combined presence within the LLM activations of the later layers of the LLM. For example, if QTR represents covered query term ratio, STF represents mean of stream length normalized term frequency and VTFIDF represents variance of tf*idf, we find high scores on average for all of QTR+STF, QTR+VTFIDF, STF+VTFIDF, QTR+STF+VTFIDF, QTR*STF, STF*VTFIDF, QTR*VTFIDF, and QTR*STF*VTFIDF when probing the last layer of the LLM. This potentially indicates that the LLM has over the course of layers learned some representation of (QTR+STF+VTFIDF)k. In Figure 6, we show the probing performance of a sum of those three features and their exponents, relative to BM25 in RankLlama2-7b and 13b models. We observe that the sum of this feature group, its square, as well as its cube consistently show a strong correlation to neuron activations within RankLlama. This analysis provides a foundation for future work to uncover and investigate complex groups of features encoded within LLM activations. By highlighting how various ranking architectures capture and represent different IR features, it opens avenues for deeper exploration into latent feature interactions and the mechanistic underpinnings of ranking-focused language models.
@@ 5. Discussion
Comparing Notes with Previous Probing Work: A number of studies have probed BERT and T5 with the aim of understanding if they encode concepts like term frequency and inverse document frequency. Formal et al. (Formal et al., 2021) study the matching process of ColBERT (Khattab and Zaharia, 2020) and conclude that the model is able to capture a notion of term importance and relies on exact matches for important terms. This is in agreement with our findings, where we find strong correlations to term matching features like covered query term number/ratio and stream length normalized term frequency within LLM activations. In a later work, Formal et al. (Formal et al., 2022) conduct a study to measure the out-of-domain zero-shot capabilities of BERT/T5 models in lexical matching on the BEIR dataset, and show that these models fail to generalize lexical matching in out-of-domain datasets or terms not seen at training time. This finding continues for RankLlama as well, where the model is unable to generalize lexical matching to terms not seen beforehand (Figure 5, RankLlama 13b OOD vs In-distribution probing discrepencies).
Different types of neurons: A monosemantic neuron is a neuron within a neural network that responds to a single, specific feature or concept in the input data (Bau et al., 2020). However, in large neural networks the networks often try to extract more features than the number of neurons. In such scenarios, the model has to perform a superposition of features on single neurons in order to compress desired features into the limited number of neurons. This gives rise to polysemantic neurons – i.e., neurons encoding more than one feature. This phenomenon is usually witnessed in the early layers of the LLM architecture and is difficult to disentangle using linear probes. There is unfortunately no known method to identify monosemantic vs polysemantic neurons via probing. Another common phenomena in sparse LLMs is when multiple neurons within a layer together represent a particular human-known context. Among the features probed in our study, $23\%$ were found in a single neuron, $32\%$ were distributed between 2-3 neurons and the other $45\%$ were distributed in 4 or more neurons.
@@
Validating Probing Results While probing techniques are widely used to understand the internal workings of LLMs, there are several limitations of using probing for this purpose. (1) Probing techniques reveal correlations between specific features and neuron activations, but not a causal analysis of the decision making process. (2) The insights gained from probing are heavily dependent on the dataset used for probing. If the dataset is not representative of the model’s typical input, the results may not accurately reflect the model’s general behavior. (3) LLMs often rely on complex interactions between multiple features. Probing techniques may fail to capture these interactions, leading to an incomplete understanding of how the model makes decisions. (4) Probing can identify features but may mislead regarding their role or significance in the model’s decision-making process. However, many of the limitations of probing can be addressed by employing various techniques, such as creating balanced probing datasets and independently validating probing results through methods like ablation studies and feature attribution analyses.
@@
To validate our probing results, we analyzed neurons in the final layer of RankLLaMA models that exhibited strong correlations ($R^{2}>0.85$) with probed features, as this layer directly influences ranking decisions. Using Shapley values  (Datta et al., 2016; Lundberg and Lee, 2017; Chowdhury et al., 2024) for feature attributions we assessed the contributions of individual neurons and neuron groups to the ranking predictions, leveraging a value function based on changes in NDCG scores for MS MARCO query-document pairs. Probing validation was conducted on 100 query-document pairs for each RankLLaMA model, and we computed average attributions across neuron groups of varying sizes. Results showed that identified neuron groups ranked within the 95th percentile in 79 out of 100 cases for RankLLaMA 7B and 84 out of 100 for RankLLaMA 13B, confirming that these neurons are instrumental in ranking decisions and validating the location of the probes.
@@ 6. Conclusion and Future Work
In this study, we probed four different large language models, each fine-tuned for passage reranking, to look for a subset of features from the MSLR dataset. These human-engineered features, which include query-only, document-only, and query-document interaction features, are recognized as significant in ranking tasks. Using a layer-wise probe, we discovered that the activations in most layers could accurately replicate select features such as Min TF*IDF, covered QT Number/Ratio, Min TF, Min(TF/L), and Mean(TF/L). This suggests that the fine-tuned Llama network deems these features important. Conversely, we found no correspondence for features like sum and max of tf*idf and sum and max of tf in the probed neurons. This indicates the absence of these features in RankLlama’s neural feature extractors. We also reported observations on feature groups that show a strong correlation with RankLlama neurons and hypothesized that some abstract features are mined by the LLM. We also observed that different LLM architectures, when fine-tuned on the same dataset with the same loss function, exhibit similar latent features when probed. Finally, we compared later features encoded between in-distribution and out-of-distribution datasets and encountered a case of the LLM’s apparently overfitting during fine-tuning. These findings enhanced our understanding of ranking LLMs and gave us generalizable insights for the design of more effective and transparent ranking models. The long-term objectives of this endeavor include: (i) identifying potential modifications to existing MSLR features, such that they further alight with LLM activations, (ii) deciphering and reverse-engineering segments of the LLM that do not correspond to recognized MSLR features, and (iii) ultimately cataloging all features deemed significant by LLMs and plugging them back into simpler statistical models to enhance the performance and interpretability of statistical ranking models. These directions will help further explain the underlying mechanisms of ranking LLMs and improve human trust into them.

### v3

@@ 4.2. Comparing Ranking LLMs
We next explore whether different ranking LLMs encode similar or distinct Information Retrieval features. To investigate this, we extend our analysis beyond the RankLlama2-7b model and compare it to our two additional fine-tuned LLMs, RankLlama3-8b and RankPythia-6.9b, as well as the RankLlama2-13b model. We fine-tune these models under an identical passage reranking framework, with an identical dataset and loss function, to ensure a fair and consistent comparison.
@@
All four ranking LLMs demonstrate broadly comparable patterns for most features. They appear closely aligned on coverage-based features—such as Covered Query Term Number and Ratio—with each scoring in the high 0.95–0.98 range. By contrast, the greatest differences emerge in TF-based statistics like Max of Term Frequency and Variance of TF/TF*IDF, where RankPythia tends to exhibit slightly higher peaks (e.g., a larger $R^{2}$ for Max of TF) than the RankLLaMA models. Additionally, all LLMs show near zero performance in encoding Sum of Stream Length Normalized TF, suggesting similarity in how these models handle term-frequency weighting. Overall, the four models encode IR features in a largely similar manner, indicating that our probing findings are generalizable and not LLM-specific.
@@ 4.3. Out-of-distribution queries and documents
We compared the probing performance of in-distribution and out-of-distribution query-document pairs in RankLlama3-8b and RankPythia-6.9b, observing similar performance for both types of inputs, consistent with the results observed for RankLlama2-7b in Figure 5. The graphs depicting these probing results are not included due to space constraints.
@@ 4.4. Feature Groups within RankLlama
In the previous subsection, we found that several MSLR features (covered QT Number, covered QT Ratio, mean(TF/L) and variance of TF*IDF) appear to be modeled within LLM activations, especially within the later layers. These features also seem to be a fair choice to model relevance based on intuition. However, these features might not be present individually but in combination with one another or individually with different exponents. To test for this possibility, we probe for combinations of these features within different layers of the LLM and find various combinations of covered query term ratio, mean of stream length normalized term frequency and variance of tf*idf. We find strong indication for their combined presence within the LLM activations of the later layers of the LLM. For example, if QTR represents covered query term ratio, STF represents mean of stream length normalized term frequency and VTFIDF represents variance of tf*idf, we find high scores on average for all of QTR+STF, QTR+VTFIDF, STF+VTFIDF, QTR+STF+VTFIDF, QTR*STF, STF*VTFIDF, QTR*VTFIDF, and QTR*STF*VTFIDF when probing the last layer of the LLM. This potentially indicates that the LLM has over the course of layers learned some representation of (QTR+STF+VTFIDF)k. In Figure 6, we show the probing performance of a sum of those three features and their exponents, relative to BM25 in RankLlama2-7b and 13b models. We observe that the sum of this feature group, its square, as well as its cube consistently show a strong correlation to neuron activations within RankLlama. This analysis provides a foundation for future work to uncover and investigate complex groups of features encoded within LLM activations. By highlighting how various ranking architectures capture and represent different IR features, it opens avenues for deeper exploration into latent feature interactions and the mechanistic underpinnings of ranking-focused language models.
@@ 5. Discussion
Validating Probing Results While probing techniques are widely used to understand the internal workings of LLMs, there are several limitations of using probing for this purpose (Belinkov, 2022). (1) Probing techniques reveal correlations between specific features and neuron activations, but not a causal analysis of the decision making process. (2) The insights gained from probing are heavily dependent on the dataset used for probing. If the dataset is not representative of the model’s typical input, the results may not accurately reflect the model’s general behavior. (3) LLMs often rely on complex interactions between multiple features. Probing techniques may fail to capture these interactions, leading to an incomplete understanding of how the model makes decisions. (4) Probing can identify features but may mislead regarding their role or significance in the model’s decision-making process. However, many of the limitations of probing can be addressed by employing various techniques, such as creating balanced probing datasets and independently validating probing results through methods like ablation studies and feature attribution analyses.
@@
To partially validate our probing results, we analyzed neurons in the final layer of RankLLaMA models that exhibited strong correlations ($R^{2}>0.85$) with probed features, as this layer directly influences ranking decisions. Approximating Shapley values  (Datta et al., 2016; Lundberg and Lee, 2017; Chowdhury et al., 2024) for feature attributions via Integrated Gradients (Sundararajan et al., 2017), we assessed the contributions of individual neurons and neuron groups to the ranking predictions, leveraging a value function based on changes in NDCG scores for MS MARCO query-document pairs. Probing validation was conducted on 100 query-document pairs for each RankLLaMA model, and we computed average attributions across neuron groups of varying sizes. Results showed that identified neuron groups ranked within the 95th percentile in 79 out of 100 cases for RankLLaMA 7B and 84 out of 100 for RankLLaMA 13B, confirming that these neurons are instrumental in ranking decisions and validating the location of the probes.
Comparing Notes with Previous Probing Work: A number of studies have probed BERT and T5 with the aim of understanding if they encode concepts like term frequency and inverse document frequency. Formal et al. (Formal et al., 2021) study the matching process of ColBERT (Khattab and Zaharia, 2020) and conclude that the model is able to capture a notion of term importance and relies on exact matches for important terms. This is in agreement with our findings, where we find strong correlations to term matching features like covered query term number/ratio and stream length normalized term frequency within LLM activations. In a later work, Formal et al. (Formal et al., 2022) conduct another study to measure the out-of-domain zero-shot capabilities of BERT/T5 models in lexical matching on the BEIR dataset, and show that these models fail to generalize lexical matching in out-of-domain datasets or terms not seen at training time. This finding continues for RankLlama as well, where the model is unable to generalize lexical matching to terms not seen beforehand (Figure 5, RankLlama 13b OOD vs In-distribution probing discrepencies).
@@ 6. Conclusion and Future Work
In this study, we probed four different large language models, each fine-tuned for passage reranking, to look for a subset of features from the MSLR dataset. These human-engineered features, which include query-only, document-only, and query-document interaction features, are recognized as significant in ranking tasks. Using a layer-wise probe, we discovered that the activations in most layers could accurately replicate select features such as covered QT Number/Ratio, Var(TF/IDF) and Mean(TF/L). This suggests that the fine-tuned Llama network deems these features important. Conversely, we found no correspondence for features like sum and max of tf*idf and sum and max of tf in the probed neurons. This indicates the absence of these features in RankLlama’s neural feature extractors. We also reported observations on feature groups that show a strong correlation with RankLlama neurons and hypothesized that some abstract features are mined by the LLM. We also observed that different LLM architectures, when fine-tuned on the same dataset with the same loss function, exhibit similar latent features when probed. Finally, we compared later features encoded between in-distribution and out-of-distribution datasets and encountered a case of the LLM’s apparently overfitting during fine-tuning. These findings enhanced our understanding of ranking LLMs and gave us generalizable insights for the design of more effective and transparent ranking models. The long-term objectives of this endeavor include: (i) identifying potential modifications to existing MSLR features, such that they further align with LLM activations, (ii) deciphering and reverse-engineering segments of the LLM that do not correspond to recognized MSLR features, and (iii) ultimately cataloging all features deemed significant by LLMs and plugging them back into simpler statistical models to enhance the performance and interpretability of statistical ranking models. These directions will help further explain the underlying mechanisms of ranking LLMs and improve human trust in them.